I can't get an answer to
1 divided by the absolute value of 1 minus 7x that is equal to 6
Its a fraction that is set equal to 6.
also, i've tried multiplying by 1 minus 7x, and get x equals 1 divided by 7 but this answer is wrong.

Answer by stanbon(57347)

1 divided by the absolute value of 1 minus 7x that is equal to 6
-----------------------------
1/[|1-7x| = 6
---
Invert both sides to get:
|1-7x| = 1/6
1-7x = 1/6 or 1-7x = -1/6
7x = (5/6) or 7x = 7/6
x = 5/42 or x = 1/6
-----------------------
Cheers,
Stan H.